OpenStudy (anonymous):

The average speed of one airplane is 200 mph faster than that of an automobile. If the automobile starts at 3 a.m. and the airplane starts from the same place at 11 a.m., they will have traveled the same number of miles by 1:00 p.m. Find the rate of the automobile.

OpenStudy (anonymous):

(x+200) * 2 = x*10

OpenStudy (anonymous):

solve for x

OpenStudy (anonymous):

Time between 11 am and 1 pm is 2 h. During this time the plane has travelled the same distance as the car has travelled in 10 h. If v is the velocity of the car The plane travels 200 * 2 + v*2 The car travels v*10. they travel the same distance 400 + 2v = 10v 400 = 10v - 2v 400 = 8v v = 50 mph and speed of plane is a very slow 250 mph.

OpenStudy (anonymous):

@sauravshakya is correct

OpenStudy (anonymous):

very correct

OpenStudy (anonymous):

A druggist has two mixtures of a certain chemical and water, one containing 10% of the chemical and the other containing 40% of the chemical. How much of each mixture should he use to make 10 ounces that contain 25% of the chemical? If x is the amount of 10% mixture used, then which expression represents the ounces of chemical contained in it?

OpenStudy (anonymous):

x ounces of 10% soln mixed with (10 – x) ounces of 40% soln. as total final volume is 10 ounces amount of chemical in x ounces = 0.1x amount of chemical in (10 – x) ounces = 0.4(10 – x) = 4 – 0.4x amount of chemical in 10 ounces = 10 × 0.25 = 2.5 chemical balance 0.1x + 4 – 0.4x = 2.5 0.3x = 1.5 x = 5 ounces of 10% mixed with (10 – 5) = 5 ounces of 40% to give 10 ounces of 25% daughter takes x hr in 1 hr 1/x work done

OpenStudy (anonymous):

i need help @sauravshakya

OpenStudy (anonymous):

is correct my answers

OpenStudy (anonymous):

PLZ wait a minute

OpenStudy (anonymous):

a;0.1x b;x c;10x i think my answer is a or c

OpenStudy (anonymous):

help please

OpenStudy (anonymous):

help